# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ 5.2 Dynamics on the regular boundary
We refer to [Sag14] for background on hyperplanes and halfspaces in $\mathrm{CAT}(0)$ cube complexes. Let $X$ be a $\mathrm{CAT}(0)$ cube complex. The Roller compactification of $X$ is defined in [Rol98, ${}^{+}$09] as follows. Let $\mathrm{HS}$ be the set of all halfspaces in $X$ (i.e. complementary components of hyperplanes). Let $\varphi:VX\to\{0,1\}^{\mathrm{HS}}$ be the map such that $\varphi(v)(h)=1$ if $v\in h$, and $\varphi(v)(h)=0$ otherwise. The Roller compactification, denoted $\overline{X}^{R}$, is the closure of $\varphi(VX)$ in $\{0,1\}^{\mathrm{HS}}$, in the topology of pointwise convergence. Thus, every point $\xi\in\overline{X}^{R}$ is a function from $\mathrm{HS}$ to $\{0,1\}$; we denote by $\langle\xi,h\rangle$ the value $\xi(h)$, and say that $h$ points towards $\xi$ if $\langle\xi,h\rangle=1$. The Roller boundary is $\partial_{R}X=\overline{X}^{R}\setminus VX$.
@@
For essentiality, notice that every hyperplane crosses an edge $e$ joining a rank 00 vertex to a rank $1$ vertex. It is enough to prove that $e$ is contained in a periodic bi-infinite geodesic (for the $\mathrm{CAT}(0)$ metric on $\mathbb{B}$).
Let $\Gamma^{c}$ be the complement graph of $\Gamma$, i.e. $V\Gamma^{c}=V\Gamma$, and two vertices of $\Gamma^{c}$ are adjacent if and only if they are non-adjacent in $\Gamma$. As $\Gamma$ is not a join, $\Gamma^{c}$ is connected. Let $a_{1}$ be the label of $e$, and let $\gamma$ be an immersed loop in $\Gamma^{c}$ based at $a_{1}$. Let $a_{1},\dots,a_{n},a_{n+1}=a_{1}$ be the vertices encountered along $\gamma$, and let $w=a_{1}\dots a_{n}$, an element of $G$. Recall that each vertex of $\mathbb{B}$ is represented by a left coset of a standard flat in $G$. We consider the following sequence of vertices in $\mathbb{B}$, alternating between rank 0 and rank 1 vertices:
@@
Consecutive members in this sequence are adjacent vertices in $\mathbb{B}$, so the above gives an edge path $Q$. We claim that $Q$ is a geodesic segment: indeed, the angle (in the sense of e.g. [BH99, Chapter II.3]) between two adjacent edges at a rank 1 vertex is $\pi$ (using that the link of any rank $1$ vertex has a bipartition into rank 00 and higher-rank vertices). And the angle between two adjacent edges at a rank 0 vertex is $\pi$ because $a_{i}$ and $a_{i+1}$ are not adjacent in $\Gamma$. This is enough to ensure that $Q$ is a geodesic segment as local geodesics in a $\mathrm{CAT}(0)$ space are global geodesics [BH99, Proposition II.1.4(2)]. Likewise, as $a_{1}$ and $a_{n}$ are not adjacent in $\Gamma$, the concatenation $\cup_{i\in\mathbb{Z}}w^{i}Q$ is a geodesic line $\ell_{w}$ in $\mathbb{B}$, which is an axis for $w$. It has a translate passing through $e$, which is the desired bi-infinite geodesic. ∎
@@ 6 Action on the right-angled building with amenable stabilizers
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $\mathbb{B}$ be its right-angled building, and let $H$ be a countable group. Assume that $G$ and $H$ are measure equivalent. Then $H$ acts on $\mathbb{B}$ with amenable vertex stabilizers. If in addition $H$ has bounded torsion, then this action can be chosen to be cocompact.
@@ 6.1 A finite-index extension of $G$ with the same transitivity as $\mathrm{Aut}(\mathbb{B})$
Throughout the section, we let $G=G_{\Gamma}$ be a right-angled Artin group. The (finite) automorphism group $\mathrm{Aut}(\Gamma)$ naturally acts on $G$ by (outer) automorphisms. We let $\hat{G}=G\rtimes\mathrm{Aut}(\Gamma)$. The action of $\mathrm{Aut}(\Gamma)$ on $G$ sends standard flat to standard flat. Therefore, the $G$-action on its right-angled building $\mathbb{B}$ extends to an action of $\hat{G}$ by cubical automorphisms, where an element $(g,\theta)\in\hat{G}$ sends a vertex representing a coset $hG_{\Lambda}$, where $\Lambda\subseteq\Gamma$ is a complete subgraph, to the vertex representing $g\theta(hG_{\Lambda})$. The importance of the group $\hat{G}$ for us is that it acts on $\mathbb{B}$ with the same transitivity as the full automorphism group $\mathrm{Aut}(\mathbb{B})$, as shown by the following lemma. This property will be crucial in order to apply Corollary 4.7.
@@
It suffices to show that if $he=e^{\prime}$ for some $h\in\mathrm{Aut}(\mathbb{B})$, then there exists $g\in\hat{G}$ such that $ge=e^{\prime}$. Recall that vertices of $\mathbb{B}$ are in one-to-one correspondence with standard flats in $G$, and elements in $\mathrm{Aut}(\mathbb{B})$ can be viewed as flat-preserving bijections of $G$. Edges of $\mathbb{B}$ are in one-to-one correspondence with codimension one inclusions of standard flats $F_{1}\subseteq F_{2}$. Let $F_{1}\subseteq F_{2}$ and $F^{\prime}_{1}\subseteq F^{\prime}_{2}$ be the inclusions of standard flats associated to the edges $e$ and $e^{\prime}$, respectively. Since every automorphism of $\mathbb{B}$ preserves the ranks of vertices (Lemma 2.5), we have $h(F_{1},F_{2})=(F^{\prime}_{1},F^{\prime}_{2})$. Let $x\in F^{\prime}_{1}$ be a vertex. Then there is a unique standard flat $\tilde{F}_{1}$ containing $x$ such that $F_{1}$ and $\tilde{F}_{1}$ have the same type. Likewise there exists a unique standard flat $\tilde{F}_{2}$ containing $\tilde{F}_{1}$ which has the same type as $F_{2}$. Then $(F_{1},F_{2})$ and $(\tilde{F}_{1},\tilde{F}_{2})$ differ by a left translation, i.e. there exists $g_{1}\in G$ such that $g_{1}(F_{1},F_{2})=(\tilde{F}_{1},\tilde{F}_{2})$. Thus $g_{1}h^{-1}(F^{\prime}_{1},F^{\prime}_{2})=(\tilde{F}_{1},\tilde{F}_{2})$. Let $x^{\prime}=g_{1}h^{-1}(x)\in\tilde{F}_{1}$. As $x$ and $x^{\prime}$ both belong to $\tilde{F}_{1}$, there exists $g_{2}\in G$ such that $g_{2}(x^{\prime})=x$ and $g_{2}(\tilde{F}_{1},\tilde{F}_{2})=(\tilde{F}_{1},\tilde{F}_{2})$. Thus $g_{2}g_{1}h^{-1}(F^{\prime}_{1},F^{\prime}_{2})=(\tilde{F}_{1},\tilde{F}_{2})$. Also $g_{2}g_{1}h^{-1}(x)=x$ belongs to $F^{\prime}_{1}\cap\tilde{F}_{1}$. Let $v_{x}$ be the vertex in $\mathbb{B}$ associated to $x$, and let $\tilde{e}$ be the edge associated to $(\tilde{F}_{1},\tilde{F}_{2})$. Let $K_{x}$ be the union of all cubes in $\mathbb{B}$ containing $v_{x}$. Note that $\tilde{e},e^{\prime}\subset K_{x}$, and $g_{2}g_{1}h^{-1}$ stabilizes $K_{x}$, sending $e^{\prime}$ to $\tilde{e}$. Note that $\mathrm{lk}(v_{x},K_{x})$ is isomorphic to the flag completion of the defining graph $\Gamma$ of $G$, i.e. vertices of $\mathrm{lk}(v_{x},K_{x})$ are in one-to-one correspondence with vertices of $\Gamma$, and a collection of vertices in $\mathrm{lk}(v_{x},K_{x})$ span a simplex whenever the associated vertices in $\Gamma$ span a complete subgraph. So the map $(g_{2}g_{1}h^{-1})_{|K_{x}}:K_{x}\to K_{x}$, and its conjugate $(x^{-1}g_{2}g_{1}h^{-1}x)_{|K_{\mathrm{id}}}:K_{\mathrm{id}}\to K_{\mathrm{id}}$, are induced by an automorphism of $\Gamma$. This means that $x^{-1}g_{2}g_{1}h^{-1}x$ has the same action on $K_{\mathrm{id}}$ as an element of $\mathrm{Aut}(\Gamma)\subset\hat{G}$. Thus there exists $g_{3}\in\hat{G}$ such that $g_{3}(e^{\prime})=\tilde{e}$, and hence $g^{-1}_{3}g_{1}(e)=e^{\prime}$, as desired. ∎
@@
Let $h\in\mathrm{Aut}(\Gamma^{e})$ and let $\mathsf{w}=h\mathsf{v}$. We aim to prove that there exists $g\in\hat{G}$ such that $\mathsf{w}=g\mathsf{v}$. When viewed as an automorphism of $\mathbb{B}$ through the isomorphism $\mathrm{Aut}(\Gamma^{e})\to\mathrm{Aut}(\mathbb{B})$ provided by Lemma 2.6, the element $h$ sends a $\mathsf{v}$-line $\ell$ (associated to a rank $1$ vertex $v\in V\mathbb{B}$) to some $\mathsf{w}$-line $\ell^{\prime}$ (associated to a rank $1$ vertex $v^{\prime}\in V\mathbb{B}$). By Corollary 6.3, there exists $g\in\hat{G}$ with $gv=v^{\prime}$. Since $g$ sends the $\mathsf{v}$-line $\ell$ to the $\mathsf{w}$-line $\ell^{\prime}$, we have $g\mathsf{v}=\mathsf{w}$, which concludes our proof. ∎
@@ 6.2 Reduction of self couplings to the right-angled building
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, and let $\Sigma$ be a self measure equivalence coupling of $\hat{G}$. Then there exists a measurable almost $(\hat{G}\times\hat{G})$-equivariant map $\Sigma\to\mathrm{Aut}(\mathbb{B})$.
@@
Since $G$ has finite index in $\hat{G}$, the space $\Sigma$ is also a self measure equivalence coupling of $G$. Let $X_{\ell},X_{r}\subseteq\Sigma$ be respective fundamental domains for the actions of $G_{\ell}=G\times\{1\}$ and $G_{r}=\{1\}\times G$ on $\Sigma$. In view of [Kid09, Lemma 2.27], we can (and will) choose $X_{\ell},X_{r}$ so that $(G\times G)\cdot(X_{\ell}\cap X_{r})=\Sigma$ up to null sets. Let $U=X_{\ell}\cap X_{r}$. As recalled in Section 4.1, there are natural measure-preserving actions $G_{\ell}\curvearrowright X_{r}$ and $G_{r}\curvearrowright X_{\ell}$, obtained through the identifications $X_{r}\approx G_{r}\backslash\Sigma$ and $X_{\ell}\approx G_{\ell}\backslash\Sigma$. Their orbits coincide on $U$, so the two corresponding measured groupoids $G_{\ell}\ltimes X_{r}$ and $G_{r}\ltimes X_{\ell}$ restrict to isomorphic measured groupoids on $U$. We denote by $\mathcal{G}$ this common measured groupoid over $U$, which is naturally equipped (up to restricting to a conull Borel subset of $U$) with two cocycles $\rho_{\ell}:\mathcal{G}\to G_{\ell}$ and $\rho_{r}:\mathcal{G}\to G_{r}$. The first two paragraphs of the proof of [HH22a, Theorem 3.19] yield a Borel map $\theta:V\Gamma^{e}\times U\to V\Gamma^{e}$ such that for every $\mathsf{v}\in V\Gamma^{e}$, there exists a partition $U^{*}=\sqcup_{i\in I}U_{i}$ of a conull Borel subset $U^{*}\subseteq U$ into at most countably many Borel subsets satisfying the following properties:
@@
Finally, using that $G$ is normal in $\hat{G}$, Lemma 4.2 and Proposition 5.1 ensure that $\bar{\theta}$ is in fact $(\hat{G}\times\hat{G})$-equivariant, which completes our proof. ∎
@@
Proposition 6.1 follows from the combination of Lemmas 6.66.7 and Corollary 6.9 below.
@@
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $H$ be a countable group, and let $\Omega$ be a measure equivalence coupling between $\hat{G}$ and $H$. Then there exist a group homomorphism $\iota:H\to\mathrm{Aut}(\mathbb{B})$ with finite kernel and a measurable almost $(\hat{G}\times H)$-equivariant map $\theta:\Omega\to\mathrm{Aut}(\mathbb{B})$, i.e. such that for all $(g,h)\in\hat{G}\times H$ and a.e. $\omega\in\Omega$, one has $\theta((g,h)\omega)=g\theta(\omega)\iota(h)^{-1}$.
@@
Let $(\Omega,\mu)$ be a measure equivalence coupling between $\hat{G}$ and $H$. Lemma 6.6 gives a measurable group homomorphism $\iota:H\to\mathrm{Aut}(\mathbb{B})$ and a measurable almost $(\hat{G}\times H)$-equivariant map $\theta:\Omega\to\mathrm{Aut}(\mathbb{B})$. In particular $\iota$ provides an action of $H$ on $\mathbb{B}$. Let $v\in V\mathbb{B}$, let $\Omega_{v}=\theta^{-1}(\mathrm{Stab}_{\mathrm{Aut}(\mathbb{B})}(v))$, let $H_{v}=\iota^{-1}(\mathrm{Stab}_{\mathrm{Aut}(\mathbb{B})}(v))$, and let $G_{v}$ (resp. $\hat{G}_{v}$) be the stabilizer of $v$ in $G$ (resp. $\hat{G}$). The equivariance of $\theta$ ensures that $\Omega_{v}$ is invariant under the action of $\hat{G}_{v}\times H_{v}$.
@@
Given a vertex $v\in V\mathbb{B}$, we denote by $V^{0}(\mathbb{B})_{\leq v}$ the set of all rank 00 vertices of $\mathbb{B}$ that are smaller than $v$ (for the partial order on $V\mathbb{B}$ introduced in Section 2.2).
@@
Assume that $H$ has bounded torsion. Then the $H$-action on the set of rank 0 vertices of $\mathbb{B}$ has finite stabilizers and finitely many orbits. In addition, for every vertex $v\in V\mathbb{B}$, the $H_{v}$-action on $V^{0}(\mathbb{B})_{\leq v}$ has finitely many orbits.
@@
The first part follows from the first conclusion of Corollary 4.7, using that $H$ has bounded torsion and that rank 00 vertices have finite $\hat{G}$-stabilizers and all belong to the same $\hat{G}$-orbit. The second part follows from the second conclusion of Corollary 4.7, applied to $V=V^{0}(\mathbb{B})_{\leq v}$: indeed $V$ is invariant under $\mathrm{Stab}_{\mathrm{Aut}(\mathbb{B})}(v)$ because $\mathrm{Aut}(\mathbb{B})$ preserves ranks of vertices (Lemma 2.5), and the action of $G_{v}$ on $V$ is transitive. ∎
@@
Assume that $H$ has bounded torsion. Then the $H$-action on $\mathbb{B}$ is cocompact.
@@
By Lemma 6.8, $H$ acts on the set of rank 0 vertices of $\mathbb{B}$ with finitely many orbits. Recall that the action of $\mathrm{Aut}(\mathbb{B})$ on $\mathbb{B}$ preserves ranks of vertices (Lemma 2.5), hence respects the order of vertices on $\mathbb{B}$. As each vertex of rank at least $1$ is lower bounded by at least one rank 0 vertex, and there is a bound $C$ such that each rank 00 vertex is smaller than at most $C$ vertices of higher rank in $\mathbb{B}$, it follows that the action of $H$ on $\mathbb{B}$ has finitely many orbits of vertices. More generally, given any vertex $s$ of $\mathbb{B}$, there are only finitely many $k$-cells that contain $s$ as a minimal rank vertex. Thus the action of $H$ on $\mathbb{B}$ has only finitely many orbits of cells, and is therefore cocompact. ∎

### v2

@@ 5.2 Dynamics on the regular boundary
We refer to [Sag14] for background on hyperplanes and halfspaces in $\mathrm{CAT}(0)$ cube complexes. Let $X$ be a $\mathrm{CAT}(0)$ cube complex. The Roller compactification of $X$ is defined in [Rol98, BCG+09] as follows. Let $\mathrm{HS}$ be the set of all halfspaces in $X$ (i.e. complementary components of hyperplanes). Let $\varphi:VX\to\{0,1\}^{\mathrm{HS}}$ be the map such that $\varphi(v)(h)=1$ if $v\in h$, and $\varphi(v)(h)=0$ otherwise. The Roller compactification, denoted $\overline{X}^{R}$, is the closure of $\varphi(VX)$ in $\{0,1\}^{\mathrm{HS}}$, in the topology of pointwise convergence. Thus, every point $\xi\in\overline{X}^{R}$ is a function from $\mathrm{HS}$ to $\{0,1\}$; we denote by $\langle\xi,h\rangle$ the value $\xi(h)$, and say that $h$ points towards $\xi$ if $\langle\xi,h\rangle=1$. The Roller boundary is $\partial_{R}X=\overline{X}^{R}\setminus VX$.
@@
For essentiality, notice that every hyperplane crosses an edge $e$ joining a rank 0 vertex to a rank $1$ vertex. It is enough to prove that $e$ is contained in a periodic bi-infinite geodesic (for the $\mathrm{CAT}(0)$ metric on $\mathbb{B}$), as such geodesic is not contained in a bounded neighbourhood of any halfspace bounded by the hyperplane dual to $e$, and any periodic geodesic is contained in a bounded neighbourhood of any $G$-orbit.
Let $\Gamma^{c}$ be the complement graph of $\Gamma$, i.e. $V\Gamma^{c}=V\Gamma$, and two vertices of $\Gamma^{c}$ are adjacent if and only if they are non-adjacent in $\Gamma$. As $\Gamma$ is not a join, $\Gamma^{c}$ is connected. Let $a_{1}$ be the label of $e$, and let $\gamma$ be a loop in $\Gamma^{c}$ based at $a_{1}$ made of a sequence $e_{1},\dots,e_{n}$ of edges in $\Gamma^{c}$. We assume $\gamma$ has at least one edge and it is immersed, i.e. for every $i\in\{1,\dots,n\}$ (considered modulo $n$), the edge $e_{i+1}$ is not equal to the edge $e_{i}$ traversed in the opposite direction. Let $a_{1},\dots,a_{n},a_{n+1}=a_{1}$ be consecutive vertices encountered along $\gamma$, and let $w$ be the product of these generators $a_{1}a_{2}\cdots a_{n-1}a_{n}$, an element of $G$.
Figure 3: An axis of $w$.
We refer to Figure 3 for the rest of the proof. Recall that each vertex of $\mathbb{B}$ is represented by a left coset of a standard flat in $G$. We consider the following sequence of vertices in $\mathbb{B}$, alternating between rank 0 and rank 1 vertices:
@@
Consecutive members in this sequence are adjacent vertices in $\mathbb{B}$, so the above gives an edge path $Q$. We claim that $Q$ is a geodesic segment: indeed, the angle (in the sense of e.g. [BH99, Chapter II.3]) between two adjacent edges at a rank 1 vertex is $\pi$ (using that the link of any rank $1$ vertex has a bipartition into rank 0 and higher-rank vertices). And the angle between two adjacent edges at a rank 0 vertex is $\pi$ because $a_{i}$ and $a_{i+1}$ are not adjacent in $\Gamma$. This is enough to ensure that $Q$ is a geodesic segment as local geodesics in a $\mathrm{CAT}(0)$ space are global geodesics [BH99, Proposition II.1.4(2)]. Likewise, as $a_{1}$ and $a_{n}$ are not adjacent in $\Gamma$, the concatenation $\cup_{i\in\mathbb{Z}}w^{i}Q$ is a geodesic line $\ell_{w}$ in $\mathbb{B}$, which is an axis for $w$. It has a translate passing through $e$, which is the desired bi-infinite geodesic. ∎
@@ 6 Action on the right-angled building with amenable stabilizers
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $\mathbb{B}$ be its right-angled building, and let $H$ be a countable group. Assume that $G$ and $H$ are measure equivalent.
Then $H$ acts on $\mathbb{B}$ with amenable vertex stabilizers. If in addition $H$ has bounded finite subgroups, then this action can be chosen to be cocompact.
@@ 6.1 A finite-index extension of $G$ with the same transitivity as $\mathrm{Aut}(\mathbb{B})$
Throughout the section, we let $G=G_{\Gamma}$ be a right-angled Artin group. The (finite) automorphism group $\mathrm{Aut}(\Gamma)$ naturally acts on $G$ by (outer) automorphisms. We let $\hat{G}=G\rtimes\mathrm{Aut}(\Gamma)$.
The action of $\mathrm{Aut}(\Gamma)$ on $G$ sends standard flat to standard flat. Therefore, the $G$-action on its right-angled building $\mathbb{B}$ extends to an action of $\hat{G}$ by cubical automorphisms, where an element $(g,\theta)\in\hat{G}$ sends a vertex representing a coset $hG_{\Lambda}$, where $\Lambda\subseteq\Gamma$ is a complete subgraph, to the vertex representing $g\theta(hG_{\Lambda})$.
The importance of the group $\hat{G}$ for us is that it acts on $\mathbb{B}$ with the same transitivity as the full automorphism group $\mathrm{Aut}(\mathbb{B})$, as shown by the following lemma. This property will be crucial in order to apply Corollary 4.8.
@@
It suffices to show that if $he=e^{\prime}$ for some $h\in\mathrm{Aut}(\mathbb{B})$, then there exists $g\in\hat{G}$ such that $ge=e^{\prime}$. Recall that vertices of $\mathbb{B}$ are in one-to-one correspondence with standard flats in $G$, and elements in $\mathrm{Aut}(\mathbb{B})$ can be viewed as flat-preserving bijections of $G$. Edges of $\mathbb{B}$ are in one-to-one correspondence with codimension one inclusions of standard flats $F_{1}\subseteq F_{2}$. Let $F_{1}\subseteq F_{2}$ and $F^{\prime}_{1}\subseteq F^{\prime}_{2}$ be the inclusions of standard flats associated to the edges $e$ and $e^{\prime}$, respectively. Since every automorphism of $\mathbb{B}$ preserves the ranks of vertices (Lemma 2.5), we have $h(F_{1},F_{2})=(F^{\prime}_{1},F^{\prime}_{2})$.
Let $x\in F^{\prime}_{1}$ be a vertex. Then there is a unique standard flat $\tilde{F}_{1}$ containing $x$ such that $F_{1}$ and $\tilde{F}_{1}$ have the same type. Likewise there exists a unique standard flat $\tilde{F}_{2}$ containing $\tilde{F}_{1}$ which has the same type as $F_{2}$. Then $(F_{1},F_{2})$ and $(\tilde{F}_{1},\tilde{F}_{2})$ differ by a left translation, i.e. there exists $g_{1}\in G$ such that $g_{1}(F_{1},F_{2})=(\tilde{F}_{1},\tilde{F}_{2})$. Thus $g_{1}h^{-1}(F^{\prime}_{1},F^{\prime}_{2})=(\tilde{F}_{1},\tilde{F}_{2})$.
Let $x^{\prime}=g_{1}h^{-1}(x)\in\tilde{F}_{1}$. As $x$ and $x^{\prime}$ both belong to $\tilde{F}_{1}$, there exists $g_{2}\in G$ such that $g_{2}(x^{\prime})=x$ and $g_{2}(\tilde{F}_{1},\tilde{F}_{2})=(\tilde{F}_{1},\tilde{F}_{2})$. Thus $g_{2}g_{1}h^{-1}(F^{\prime}_{1},F^{\prime}_{2})=(\tilde{F}_{1},\tilde{F}_{2})$. Also $g_{2}g_{1}h^{-1}(x)=x$ belongs to $F^{\prime}_{1}\cap\tilde{F}_{1}$. Let $v_{x}$ be the vertex in $\mathbb{B}$ associated to $x$, and let $\tilde{e}$ be the edge associated to $(\tilde{F}_{1},\tilde{F}_{2})$. Let $K_{x}$ be the union of all cubes in $\mathbb{B}$ containing $v_{x}$. Note that $\tilde{e},e^{\prime}\subset K_{x}$, and $g_{2}g_{1}h^{-1}$ stabilizes $K_{x}$, sending $e^{\prime}$ to $\tilde{e}$. Note that $\mathrm{lk}(v_{x},K_{x})$ is isomorphic to the flag completion of the defining graph $\Gamma$ of $G$, i.e. vertices of $\mathrm{lk}(v_{x},K_{x})$ are in one-to-one correspondence with vertices of $\Gamma$, and a collection of vertices in $\mathrm{lk}(v_{x},K_{x})$ span a simplex whenever the associated vertices in $\Gamma$ span a complete subgraph. So the map $(g_{2}g_{1}h^{-1})_{|K_{x}}:K_{x}\to K_{x}$, and its conjugate $(x^{-1}g_{2}g_{1}h^{-1}x)_{|K_{\mathrm{id}}}:K_{\mathrm{id}}\to K_{\mathrm{id}}$, are induced by an automorphism of $\Gamma$. This means that $x^{-1}g_{2}g_{1}h^{-1}x$ has the same action on $K_{\mathrm{id}}$ as an element of $\mathrm{Aut}(\Gamma)\subset\hat{G}$. Thus there exists $g_{3}\in\hat{G}$ such that $g_{3}(e^{\prime})=\tilde{e}$, and hence $g^{-1}_{3}g_{1}(e)=e^{\prime}$, as desired. ∎
@@
Let $h\in\mathrm{Aut}(\Gamma^{e})$ and let $\mathsf{w}=h\mathsf{v}$. We aim to prove that there exists $g\in\hat{G}$ such that $\mathsf{w}=g\mathsf{v}$.
When viewed as an automorphism of $\mathbb{B}$ through the isomorphism $\mathrm{Aut}(\Gamma^{e})\to\mathrm{Aut}(\mathbb{B})$ provided by Lemma 2.6, the element $h$ sends a $\mathsf{v}$-line $\ell$ (associated to a rank $1$ vertex $v\in V\mathbb{B}$) to some $\mathsf{w}$-line $\ell^{\prime}$ (associated to a rank $1$ vertex $v^{\prime}\in V\mathbb{B}$). By Corollary 6.3, there exists $g\in\hat{G}$ with $gv=v^{\prime}$. Since $g$ sends the $\mathsf{v}$-line $\ell$ to the $\mathsf{w}$-line $\ell^{\prime}$, we have $g\mathsf{v}=\mathsf{w}$, which concludes our proof. ∎
@@ 6.2 Reduction of self couplings to the right-angled building
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, and let $\Sigma$ be a self measure equivalence coupling of $\hat{G}$.
Then there exists a measurable almost (G^×G^)($\hat$$\times\hat{G}$)-equivariant map $\Sigma\to\mathrm{Aut}(\mathbb{B})$.
@@
Since $G$ has finite index in $\hat{G}$, the space $\Sigma$ is also a self measure equivalence coupling of $G$.
Let $X_{\ell},X_{r}\subseteq\Sigma$ be respective fundamental domains for the actions of $G_{\ell}=G\times\{1\}$ and $G_{r}=\{1\}\times G$ on $\Sigma$. In view of [Kid09, Lemma 2.27], we can (and will) choose $X_{\ell},X_{r}$ so that $(G\times G)\cdot(X_{\ell}\cap X_{r})=\Sigma$ up to null sets. Let $U=X_{\ell}\cap X_{r}$.
As recalled in Section 4.1, there are natural measure-preserving actions $G_{\ell}\curvearrowright X_{r}$ and $G_{r}\curvearrowright X_{\ell}$, obtained through the identifications $X_{r}\approx G_{r}\backslash\Sigma$ and $X_{\ell}\approx G_{\ell}\backslash\Sigma$. Their orbits coincide on $U$, so the two corresponding measured groupoids $G_{\ell}\ltimes X_{r}$ and $G_{r}\ltimes X_{\ell}$ restrict to isomorphic measured groupoids on $U$. We denote by $\mathcal{G}$ this common measured groupoid over $U$, which is naturally equipped (up to restricting to a conull Borel subset of $U$) with two cocycles $\rho_{\ell}:\mathcal{G}\to G_{\ell}$ and $\rho_{r}:\mathcal{G}\to G_{r}$.
The first two paragraphs of the proof of [HH22a, Theorem 3.19] yield a Borel map $\theta:V\Gamma^{e}\times U\to V\Gamma^{e}$ such that for every $\mathsf{v}\in V\Gamma^{e}$, there exists a partition $U^{*}=\sqcup_{i\in I}U_{i}$ of a conull Borel subset $U^{*}\subseteq U$ into at most countably many Borel subsets satisfying the following properties:
@@
Finally, using that $G$ is normal in $\hat{G}$, Lemma 4.3 and Proposition 5.1 ensure that $\bar{\theta}$ is in fact (G^×G^)($\hat$$\times\hat{G}$)-equivariant, which completes our proof. ∎
@@
Proposition 6.1 follows from the combination of Lemmas 6.6, 6.7 and Corollary 6.9 below.
@@
Let $G$ be a right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, let $H$ be a countable group, and let $\Omega$ be a measure equivalence coupling between $\hat{G}$ and $H$.
Then there exist a group homomorphism $\iota:H\to\mathrm{Aut}(\mathbb{B})$ with finite kernel and a measurable almost $(\hat{G}\times H)$-equivariant map $\theta:\Omega\to\mathrm{Aut}(\mathbb{B})$, i.e. such that for all $(g,h)\in\hat{G}\times H$ and a.e. $\omega\in\Omega$, one has $\theta((g,h)\omega)=g\theta(\omega)\iota(h)^{-1}$.
@@
Let $(\Omega,\mu)$ be a measure equivalence coupling between $\hat{G}$ and $H$. Lemma 6.6 gives a measurable group homomorphism $\iota:H\to\mathrm{Aut}(\mathbb{B})$ and a measurable almost $(\hat{G}\times H)$-equivariant map $\theta:\Omega\to\mathrm{Aut}(\mathbb{B})$. In particular $\iota$ provides an action of $H$ on $\mathbb{B}$.
Let $v\in V\mathbb{B}$, let $\Omega_{v}=\theta^{-1}(\mathrm{Stab}_{\mathrm{Aut}(\mathbb{B})}(v))$, let $H_{v}=\iota^{-1}(\mathrm{Stab}_{\mathrm{Aut}(\mathbb{B})}(v))$, and let $G_{v}$ (resp. $\hat{G}_{v}$) be the stabilizer of $v$ in $G$ (resp. $\hat{G}$). The equivariance of $\theta$ ensures that $\Omega_{v}$ is invariant under the action of $\hat{G}_{v}\times H_{v}$.
@@
Given a vertex $v\in V\mathbb{B}$, we denote by $V^{0}(\mathbb{B})_{\leq v}$ the set of all rank 0 vertices of $\mathbb{B}$ that are smaller than $v$ (for the partial order on $V\mathbb{B}$ introduced in Section 2.2).
@@
Assume that $H$ has bounded finite subgroups. Then the $H$-action on the set of rank 0 vertices of $\mathbb{B}$ has finite stabilizers and finitely many orbits. In addition, for every vertex $v\in V\mathbb{B}$, the $H_{v}$-action on $V^{0}(\mathbb{B})_{\leq v}$ has finitely many orbits.
@@
The first part follows from the first conclusion of Corollary 4.8, using that $H$ has bounded finite subgroups and that rank 0 vertices have finite $\hat{G}-$stabilizers and all belong to the same $\hat{G}-$orbit. The second part follows from the second conclusion of Corollary 4.8, applied to $V=V^{0}(\mathbb{B})_{\leq v}$: indeed $V$ is invariant under $\mathrm{Stab}_{\mathrm{Aut}(\mathbb{B})}(v)$ because $\mathrm{Aut}(\mathbb{B})$ preserves ranks of vertices (Lemma 2.5), and the action of $G_{v}$ on $V$ is transitive. ∎
@@
Assume that $H$ has bounded finite subgroups. Then the $H$-action on $\mathbb{B}$ is cocompact.
@@
By Lemma 6.8, $H$ acts on the set of rank 0 vertices of $\mathbb{B}$ with finitely many orbits. Recall that the action of $\mathrm{Aut}(\mathbb{B})$ on $\mathbb{B}$ preserves ranks of vertices (Lemma 2.5), hence respects the order of vertices on $\mathbb{B}$. As each vertex of rank at least $1$ is lower bounded by at least one rank 0 vertex, and there is a bound $C$ such that each rank 0 vertex is smaller than at most $C$ vertices of higher rank in $\mathbb{B}$, it follows that the action of $H$ on $\mathbb{B}$ has finitely many orbits of vertices. More generally, given any vertex $s$ of $\mathbb{B}$, there are only finitely many $k$-cells that contain $s$ as a minimal rank vertex. Thus the action of $H$ on $\mathbb{B}$ has only finitely many orbits of cells, and is therefore cocompact. ∎